OpenStudy (anonymous):

What is the sum of the measures of the interior angles of a regular polygon where a single exterior angle measures 72°?

OpenStudy (anonymous):

more help please? I know the formula but how do i find out how many sides there are for n?

OpenStudy (anonymous):

is it 252?

OpenStudy (anonymous):

I did 180n - 72n = 108 and then 360 - 108 = 252?

OpenStudy (anonymous):

Wait, I'm confused now too...@campbell_st care to clarify?

OpenStudy (anonymous):

To solve for N i just did 108n - 72n = 108 .. then 360 - 108 = 252?

OpenStudy (campbell_st):

the sum of the exterior angles of a polygon is 360 there are 2 parts, the number of sides( or angles is found by) Number of sides(angles) n = 360/72 then the interior angle sum is Angle Sum = (n - 2)*180

OpenStudy (anonymous):

Oh. I see...I misread the question :/

OpenStudy (campbell_st):

ooops n = number of sides

OpenStudy (anonymous):

i thought so. so how do i solve for N first?

OpenStudy (anonymous):

Yup. Following what ^ said, 360/72 = n. If so, what is n?

OpenStudy (campbell_st):

yes thats correct,,,,

OpenStudy (anonymous):

So N = 5

OpenStudy (anonymous):

Yes. Now, plug it in here: \[(n - 2)180\]

OpenStudy (anonymous):

540?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

ah, thankyou ! :D

OpenStudy (anonymous):

np :)